find the principal which
amount at simple interest to #600 in 2 years at
5% per annum.

Answer by stanbon(75887)

find the principal which
amount at simple interest to #600 in 2 years at
5% per annum.
------------------
A(t) = P*r*t
----
600 = P*0.05*2
-----
P = 600/(0.1)
P = $6000.00
===============
Cheers,
Stan H.
------------------